AIEEE 2005 : if we consider that 1/6 , in place of 1/12 , mass of carbon atom is taken to be the relative atomic mass unit , the mass of one mole of a substance : a) decrease twice b) increase two fold c) remain unchanged d) be a function of the molecular mass of the substance . The answer is remains unchanged . Why is it so ? Will all the other terms like atomic mass , molecular mass also remain unchanged ? Wont the relative atomic mass and atomic mass become halved ? Pls explain

Grade:12th Pass

2 Answers

Sunil Kumar FP
10 years ago
all will remain unchanged because mass of 1 mole remians unchanged it mean s atomic mass or molecupar mass of the element or compound.
sloution:-
mass of 1 mole=Na× mss of 1 atom
when iit is 1/6 then amu^ =2 amu
Na will become Na^=Na/2 now 6 gm wiol corresponds to avagadro no that is Na/2
so mulriplying=Na/2×2amu =A or M(atomic or molecular mass)

thanks and regards
sunil kr
askIItian faculty

SANIKA
10 years ago
Why does mass of one mole remain unchanged?